If someone 50% of their cereal 20% of their fruit 15% of their milk and 15% of their juice how much did they eat total
Answer by dkppathak(439)

50% of their cereal =50/100=1/2 part LEFT =1/2 part
20% of their fruit =20/100 =1/5 part LEFT=4/5 part
15% of their milk =15/100 =3/20 part LEFT=17/20 part
and
15% of their juice =15/100=3/20 part LEFT=17/20 part
how much did they eat total
total eat=1/2+1/5+3/20+3/20 by LCM LEFT =1/2+4/5+17/20+17/20 part
=10+4+3+3/20
=20/20=1 PART CONSUMED LEFT PART =10+16+17+17/20
=60/20 part
CONSUMED PART=1 LEFT PART =3 PART

TOTAL EAT PART =1/4 LEFT PART =3/4
Answer